# Anonymization Prompt Learning for Facial Privacy-Preserving Text-to-Image Generation

### v2

@@ 4.3 Quality Preservation
We strive to ensure our anonymization is natural and specific to identifiable facial images. When requested to generate specific individuals, models equipped with the Anonymization Prompts should return realistic facial images with false identities, but retain facial attributes similar to those of the requested identity, as outlined in section 3.4. In addition, this modification should not interfere with generation qualities of unrelated images, i.e., concatenating the Anonymization Prompt to prompts of non-facial images and non-identity-specific facial images should not significantly change the outputs.
In this section, we comprehensively evaluate the ability of our method to meet these requirements. We first study the quality of 4000 facial images generated with identity-specific prompts from our dataset $S_{test}$. We calculate the FID [30] score of our generated faces against those generated by the original model. For attribute evaluation, we calculate the Attribute Accuracy (Attr-ACC) of generated facial images, which is the average cosine similarity between the facial attribute features [33] of the generated faces and the ground truths. For non-identity-specific generations, we generate $30k$ images with prompts from COCO-30k [31]. We evaluate FID scores between images generated by modified models and the original Stable Diffusion [1], and evaluate CLIP scores by averaging the distances between all generated images and their corresponding text prompts.
A direct baseline of our task is to apply concept editing methods [8, 9, 7, 16] and remove facial identities as target concepts. We compare our results with state-of-the-art concept editing, namely ESD [8], Concept Ablation (CA) [9], UCE [16] and Forget-me-not (FMN) [7]. We run their open training code on SD-v1.5 [1] but substitute their editing concepts with identities in our training set.
Table 2: Quantitative results of quality preservation. We report performance on ID-specific generation and non-ID-specific generation, which are respectively evaluated with ID-specific prompts and prompts from COCO-30k [31]. Best results are in bold and second bests are underlined.
@@
Results in table 2 demonstrate that our method best satisfies all requirements for a successful anonymization. It not only generates high quality facial images with low identity accuracies, but also preserves facial attribute information of the requested identities. On generating COCO [31] images, APL achieves the highest image quality and text fidelity, respectively quantified by FID and CLIP scores, surpassing all compared methods.
In contrast, the compared methods fail to meet the demand of our anonymization task. After training with our dataset, some methods exhibit difficulties in generating detectable faces, therefore we do not report their ID-ACC and Attr-ACC. For example, FMN [7] and UCE [16] produce high FID scores when prompted with specific identities, such that a large number of images do not contain any faces. Simultaneously, they generate COCO images with high FID scores and low CLIP scores, as the resulting images are neither realistic nor related to the input prompts. This is outcome is consistent with existing research [16], which suggests that editing too many concepts with existing methods may significantly impact the generation quality.
Meanwhile, we show that some methods fail to meet the requirements despite relatively maintaining generation quality. For example, models trained by ESD [8] generates random scenery images when prompted with specific identities, as shown in fig. 4. This occurs because ESD [8] aligns its target concepts to empty strings, effectively removing the concepts entirely. Similarly, Concept Ablation [9] selects anchor concepts, such as "a person", as optimization targets. Although it generates facial images with low identity accuracies, this is at the cost of minimal facial attribute similarity with the requested identity, as indicated by a markedly lower Attr-ACC. As discussed in section 3.4, we consider such anonymization excessive, as they produce completely irrelevant faces or no faces at all. We visualize excessive anonymization in fig. 4.
Figure 4: Excessive anonymization observed in compared methods. Empty string targets of ESD [8] lead to random scenery generations and category label targets of Concept Ablation [9] generate every identity as a random person. Our prompts curated with abundant attribute information strikes a balance between erasure of semantics and attaining original attributes.
@@ 4.4 Anonymizing New Identities
Personalization methods, such as Dreambooth [34], Textual Inversion [35] and HyperDreambooth [5], enable diffusion models to acquire new concepts, including highly accurate facial identities [5]. This capability significantly expands deepfake threats, affecting individuals beyond those in the model’s training dataset. Meanwhile, these methods also undermine the effectiveness of traditional concept editing methods [8, 9], as studies show that concepts removed by them can be re-introduced to the model with the same personalization procedure [6, 36].
@@
Figure 5: We evaluate the anonymization performance of our Anonymization Prompt (AP) on identites learned after APL training. (a) Identity accuracies of images generated with and without the Anonymization Prompt at different Dreambooth [34] iterations. The prompt consistently lowers identity accuracies on newly-learned identities. (b) An example of anonymizing a new identity [5], which is linked to a token [V] by Dreambooth [34]. The prompt significantly reduces the identity accuracy of the generated image.
In specific, we introduce new identities to Stable Diffusion v1.5 [1] with Dreambooth [34] and evaluate the anonymization performance of the Anonymization Prompt on these identities. We sample 50 identities from the VGGFace-2 [37] dataset, collecting high-resolution photos for each identity for personalization. We obtain a personalized Stable Diffusion for each of those identities with Dreambooth [34], so that resulting models are able to accurately generate their corresponding identities with a pre-determined token. Models are fine-tuned for 800 iterations, which is sufficient for the base model to learn accurate identities. During training, we generate images of the corresponding identities every 100 iteration, both with and without applying the Anonymization Prompt, to measure their identity accuracies. The results across all identities are then aggregated to form the curves in fig. 5(a).
Owing to the effectiveness of Dreambooth [34], both curves exhibit a clear increasing trend. However, in cases where identities are generated with the Anonymization Prompt, the identity accuracies consistently remain lower compared to those generated without the Anonymization Prompt across all iterations. With fewer iterations, the prompt effectively reduces the identity accuracies to nearly half of the base model. At 800 iteratieons, generations with the prompt achieves an average identity accuracy of 0.43, markedly lower than the baseline of 0.53, demonstrating effective anonymizations on these new identities when Dreambooth [34] is trained until convergence. We visualize the effect of the Anonymization Prompt on one learned identity in fig. 5(b). At the same training iteration of Dreambooth [34], the identity generated without the prompt closely resembles the ground truth identity, whereas the one generated with the prompt is much less recognizable, showcasing successful anonymization.
@@ 4.5 Ablation Study
Figure 6: Effects of training iterations. (a) An identity generated by RealisticVision [10] that employs the Anonymization Prompts of different training iterations. Different anonymization levels of the same identity is observed. (b) Identity accuracies computed across the test set. Users of our method may choose to use the Anonymization Prompts at different training iterations according to their own policies of content moderation.
Training iterations. We monitor the identity accuracies during the entire training process, and observe that the accuracies for all models progressively decrease over time in fig. 6(a). This effect is visualized with one identity in fig. 6(b). Both visual effects and quantitative measures indicate a steadily increasing level of anonymization. In light of these findings, text-to-image service providers may choose to employ the Anonymization Prompts at various stages of training to enforce different levels of anonymization in alignment with their content moderation policies.
@@
Regularization. We ablate the regularization loss $L_{Reg}$, which trains the Anonymization Prompt to be ineffective when concatenated to non-identity-specific prompts. We demonstrate in fig. 7 that without regularization, models generate unexpected depictions of people, which is a significant violation of semantic alignment. Since APL is a prompt learning method, training without the regularization of non-face images causes the model to link the prompt to the concept of faces, resulting in above situations. Therefore, training the prompt to ensure its ineffectiveness on non-identity-specific prompts is essential for our method to be useful in practice.
@@
Prompt length. The length of the Anonymization Prompt is the number of learnable token that constitutes our prompt prefix. We study the optimal length of the prompt by evaluating Identity Accuracies with prompts of different lengths. As shown in  table 3, we find that $m=10$ strikes a balance between effective anonymization and computational efficiency. Shorter prompts are less effective at anonymization. Longer prompts only offer marginal improvements, while requiring more computational and storage resources.
@@
Dataset Size. The Anonymization Prompt enables the anonymization of any identities, including those unseen during training and those learned through fine-tuning. We show that this generalization ability has its roots in the number of training identities. Results are presented in  fig. 8. With more training identities, we observe a consistent decrease in both average and maximum Identity Accuracies. This suggests that incorporating more identities enables the model to identify more precise directions from $c_{1}$ to $c_{2}$ in the latent space, where $c_{1}$ represents the prompts specifiying identities and $c_{2}$ are prompts encapsulating the demographic information and facial attributes of these identities. An accurate estimation of this direction guides the model to anonymize more identities successfully. We expect the identity accuracies to further drop on unseen diffusion models if larger datasets with more diverse attributes are available.
@@ Appendix A Samples from collected dataset
We present some samples from our collected dataset $S_{ID}$ in table 5. We reiterate that the attribute information is automatically generated by ChatGPT [27], and they may not accurately reflect the actual characteristics of these individuals.
Table 5: Six samples from our collected dataset $S_{ID}$. We collect a diverse set of individuals and generate their attribute description with ChatGPT [27].
@@
The collection of this dataset can be straightforwardly reproduced by querying ChatGPT [27] with a list of celebrity names, with prompts formatted as follows:

### v3

@@ 4.3 Quality Preservation
We strive to ensure our anonymization is natural and specific to identifiable facial images. When requested to generate specific individuals, models equipped with the Anonymization Prompts should return realistic facial images with false identities, but retain facial attributes similar to those of the requested identity, as outlined in section˜3.4. In addition, this modification should not interfere with generation qualities of unrelated images, i.e., concatenating the Anonymization Prompt to prompts of non-facial images and non-identity-specific facial images should not significantly change the outputs.
In this section, we comprehensively evaluate the ability of our method to meet these requirements. We first study the quality of 4000 facial images generated with identity-specific prompts from our dataset $S_{test}$. We calculate the FID [cleanfid] score of our generated faces against those generated by the original model. For attribute evaluation, we calculate the Attribute Accuracy (Attr-ACC) of generated facial images, which is the average cosine similarity between the facial attribute features [farl] of the generated faces and the ground truths. For non-identity-specific generations, we generate $30k$ images with prompts from COCO-30k [coco]. We evaluate FID scores between images generated by modified models and the original Stable Diffusion [ldm], and evaluate CLIP scores by averaging the distances between all generated images and their corresponding text prompts.
A direct baseline of our task is to apply concept editing methods [erase, ablating, forgetmenot, uce] and remove facial identities as target concepts. We compare our results with state-of-the-art concept editing, namely ESD [erase], Concept Ablation (CA) [ablating], UCE [uce] and Forget-me-not (FMN) [forgetmenot]. We run their open training code on SD-v1.5 [ldm] but substitute their editing concepts with identities in our training set.
Table 2: Quantitative results of quality preservation. We report performance on ID-specific generation and non-ID-specific generation, which are respectively evaluated with ID-specific prompts and prompts from COCO-30k [coco]. Best results are in bold and second bests are underlined.
@@
Results in table˜2 demonstrate that our method best satisfies all requirements for a successful anonymization. It not only generates high quality facial images with low identity accuracies, but also preserves facial attribute information of the requested identities. On generating COCO [coco] images, APL achieves the highest image quality and text fidelity, respectively quantified by FID and CLIP scores, surpassing all compared methods.
In contrast, the compared methods fail to meet the demand of our anonymization task. After training with our dataset, some methods exhibit difficulties in generating detectable faces, therefore we do not report their ID-ACC and Attr-ACC. For example, FMN [forgetmenot] and UCE [uce] produce high FID scores when prompted with specific identities, such that a large number of images do not contain any faces. Simultaneously, they generate COCO images with high FID scores and low CLIP scores, as the resulting images are neither realistic nor related to the input prompts. This is outcome is consistent with existing research [uce], which suggests that editing too many concepts with existing methods may significantly impact the generation quality.
Meanwhile, we show that some methods fail to meet the requirements despite relatively maintaining generation quality. For example, models trained by ESD [erase] generates random scenery images when prompted with specific identities, as shown in fig.˜4. This occurs because ESD [erase] aligns its target concepts to empty strings, effectively removing the concepts entirely. Similarly, Concept Ablation [ablating] selects anchor concepts, such as "a person", as optimization targets. Although it generates facial images with low identity accuracies, this is at the cost of minimal facial attribute similarity with the requested identity, as indicated by a markedly lower Attr-ACC. As discussed in section˜3.4, we consider such anonymization excessive, as they produce completely irrelevant faces or no faces at all. We visualize excessive anonymization in fig.˜4.
Figure 4: Excessive anonymization observed in compared methods. Empty string targets of ESD [erase] lead to random scenery generations and category label targets of Concept Ablation [ablating] generate every identity as a random person. Our prompts curated with abundant attribute information strikes a balance between erasure of semantics and attaining original attributes.
@@ 4.4 Anonymizing New Identities
Personalization methods, such as Dreambooth [dreambooth], Textual Inversion [textualinversion] and HyperDreambooth [hyperdreambooth], enable diffusion models to acquire new concepts, including highly accurate facial identities [hyperdreambooth]. This capability significantly expands deepfake threats, affecting individuals beyond those in the model’s training dataset. Meanwhile, these methods also undermine the effectiveness of traditional concept editing methods [erase, ablating], as studies show that concepts removed by them can be re-introduced to the model with the same personalization procedure [imma, ringabell].
@@
Figure 5: We evaluate the anonymization performance of our Anonymization Prompt (AP) on identites learned after APL training. (a) Identity accuracies of images generated with and without the Anonymization Prompt at different Dreambooth [dreambooth] iterations. The prompt consistently lowers identity accuracies on newly-learned identities. (b) An example of anonymizing a new identity [hyperdreambooth], which is linked to a token [V] by Dreambooth [dreambooth]. The prompt significantly reduces the identity accuracy of the generated image.
In specific, we introduce new identities to Stable Diffusion v1.5 [ldm] with Dreambooth [dreambooth] and evaluate the anonymization performance of the Anonymization Prompt on these identities. We sample 50 identities from the VGGFace-2 [vggface2] dataset, collecting high-resolution photos for each identity for personalization. We obtain a personalized Stable Diffusion for each of those identities with Dreambooth [dreambooth], so that resulting models are able to accurately generate their corresponding identities with a pre-determined token. Models are fine-tuned for 800 iterations, which is sufficient for the base model to learn accurate identities. During training, we generate images of the corresponding identities every 100 iteration, both with and without applying the Anonymization Prompt, to measure their identity accuracies. The results across all identities are then aggregated to form the curves in fig.˜5(a).
Owing to the effectiveness of Dreambooth [dreambooth], both curves exhibit a clear increasing trend. However, in cases where identities are generated with the Anonymization Prompt, the identity accuracies consistently remain lower compared to those generated without the Anonymization Prompt across all iterations. With fewer iterations, the prompt effectively reduces the identity accuracies to nearly half of the base model. At 800 iteratieons, generations with the prompt achieves an average identity accuracy of 0.43, markedly lower than the baseline of 0.53, demonstrating effective anonymizations on these new identities when Dreambooth [dreambooth] is trained until convergence. We visualize the effect of the Anonymization Prompt on one learned identity in fig.˜5(b). At the same training iteration of Dreambooth [dreambooth], the identity generated without the prompt closely resembles the ground truth identity, whereas the one generated with the prompt is much less recognizable, showcasing successful anonymization.
@@ 4.5 Ablation Study
Figure 6: Effects of training iterations. (a) An identity generated by RealisticVision [realisticvision] that employs the Anonymization Prompts of different training iterations. Different anonymization levels of the same identity is observed. (b) Identity accuracies computed across the test set. Users of our method may choose to use the Anonymization Prompts at different training iterations according to their own policies of content moderation.
Training iterations. We monitor the identity accuracies during the entire training process, and observe that the accuracies for all models progressively decrease over time in fig.˜6(a). This effect is visualized with one identity in fig.˜6(b). Both visual effects and quantitative measures indicate a steadily increasing level of anonymization. In light of these findings, text-to-image service providers may choose to employ the Anonymization Prompts at various stages of training to enforce different levels of anonymization in alignment with their content moderation policies.
@@
Regularization. We ablate the regularization loss $L_{Reg}$, which trains the Anonymization Prompt to be ineffective when concatenated to non-identity-specific prompts. We demonstrate in fig.˜7 that without regularization, models generate unexpected depictions of people, which is a significant violation of semantic alignment. Since APL is a prompt learning method, training without the regularization of non-face images causes the model to link the prompt to the concept of faces, resulting in above situations. Therefore, training the prompt to ensure its ineffectiveness on non-identity-specific prompts is essential for our method to be useful in practice.
@@
Prompt length. The length of the Anonymization Prompt is the number of learnable token that constitutes our prompt prefix. We study the optimal length of the prompt by evaluating Identity Accuracies with prompts of different lengths. As shown in  table˜3, we find that $m=10$ strikes a balance between effective anonymization and computational efficiency. Shorter prompts are less effective at anonymization. Longer prompts only offer marginal improvements, while requiring more computational and storage resources.
@@
Dataset Size. The Anonymization Prompt enables the anonymization of any identities, including those unseen during training and those learned through fine-tuning. We show that this generalization ability has its roots in the number of training identities. Results are presented in  fig.˜8. With more training identities, we observe a consistent decrease in both average and maximum Identity Accuracies. This suggests that incorporating more identities enables the model to identify more precise directions from $c_{1}$ to $c_{2}$ in the latent space, where $c_{1}$ represents the prompts specifiying identities and $c_{2}$ are prompts encapsulating the demographic information and facial attributes of these identities. An accurate estimation of this direction guides the model to anonymize more identities successfully. We expect the identity accuracies to further drop on unseen diffusion models if larger datasets with more diverse attributes are available.
@@ Appendix A Samples from collected dataset
We present some samples from our collected dataset $S_{ID}$ in table˜5. We reiterate that the attribute information is automatically generated by ChatGPT [chatgpt], and they may not accurately reflect the actual characteristics of these individuals.
Table 5: Six samples from our collected dataset $S_{ID}$. We collect a diverse set of individuals and generate their attribute description with ChatGPT [chatgpt].
@@
The collection of this dataset can be straightforwardly reproduced by querying ChatGPT [chatgpt] with a list of celebrity names, with prompts formatted as follows: